Suppose a company has 3 products. Product A has a contribution margin per unit of $15, product B has a contribution margin per unit of $25, and product C has a contribution margin per unit of $32. For every 3 units sold of product A the company sells 2units of product B and 5units of product C. The company has fixed costs of $835,000. Suppose that the company wants to make an after-taxprofit of $2,450,000 and has a tax rate of 30%, how many units of each product must the company sell to achieve this goal?
Answer #1

After tax profit = 2,450,000
Before tax profit = 2,450,000 / 70% = $ 3,500,000

Weighted average contribution = 15 X 3 / 10 + 25 X 2 / 10 + 32 X 5 / 10 = 4.5 + 5 + 16 = 25.5

Target sale = (3,500,000 + 835,000) / 25.5 = 170,000 units

A = 170,000 X 3 / 10 = 51,000 units
B = 170,000 X 2 / 10 = 34,000 units
C = 170,000 X 5 / 10 = 85,000 units